OpenStudy (anonymous):

a car travals 150km in 3 hours.What is the average speed? how far will it travel (i) five hours (ii) 45 minutes

OpenStudy (anonymous):

Avg speed in km/h=50 (i)50*5=250km (ii)50*3/4=37.5km